OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x - 4)(2x - 1)(x + 7)?

OpenStudy (lgbasallote):

equate each factor to zero x - 4 = 0 2x - 1= 0 x+7 = 0 from there solve for the three x's

OpenStudy (lgbasallote):

uhmm need more help @tien.lam12 ?

OpenStudy (anonymous):

thats it thanks

OpenStudy (lgbasallote):

owkay :)

OpenStudy (anonymous):

Set all the equation to 0...and then solve

OpenStudy (anonymous):

So you have: x - 4 = 0, which becomes x = 4 2x - 1 = 0, which becomes 2x = 1, then x = 1/2 x + 7 = 0, so x = -7 Hope this helps. Sean